The sum of two angles is 119°. One angle is 22° less than twice the other. Find the angles.


a. 68° and 51° b. 47° and 72° c. 45° and 74° d. 45° and 68°

Answer by edjones(7569)

Let the angles be x & y.
x+y=119
x+22=2y
.
x=119-y
119-y+22=2y substitute 119-y for x
140-y=2y
3y=141
y=47
x=72
.
Ed